OpenStudy (xxcookiequeenxx):

Which equations in point-slope form are equations of the line that pass through the points (−7, 5) and (−4, 1)? Choose exactly two answers that are correct. A. Y-7=-3 (X-5) - 4 B. Y-5=-4 (X+7) - 3 C. Y+4=-3 (X-1) - 4 D. y-1=-4(X+4) - 3

OpenStudy (xxcookiequeenxx):

@Cammie21

OpenStudy (anonymous):

WHAT

OpenStudy (xxcookiequeenxx):

just answer the question chill out

OpenStudy (anonymous):

I'm not good at math. I have a d- in that class

OpenStudy (texaschic101):

(1-5) / (-4- (-7) = -4/3 y - y1 = m(x - x1 B and D